Compute the determinant of :
∣−2−3∣
∣1 4∣

a. 5
b. -5
c. 4
d. -4
1
Expert's answer
2014-04-14T09:33:01-0400
The determinant of the matrix 2\times 2 can be found using the rule
| a b |
| c d | = ad-bc.
Therefore we have
∣−2 −3∣
∣1 4∣=-8+3=-5
Answer: b.